Stanley, Inc. has a new fancy product. The investment cost is expected to be $28 million and will return $8.5 million for 5 years in net cash flows. The ratio of debt to equity (D/E) is 3 to 1. The cost of equity is 14%, the pretax cost of debt is 7.5%, and the tax rate is 25%. The appropriate discount rate, assuming average risk, is:
Answer #1

Answer:

WACC = (Weight of Debt * After Tax Cost of Debt) + (Weight of Equity * Cost of Equity)

After Tax Cost of Debt = 7.50% * (1 – 0.25)
After Tax Cost of Debt = 5.625%

Weight of Debt = 3/ 4 or 75%
Weight of Equity = 1/ 4 or 25%

WACC = (0.75 * 5.625%) + (0.25 * 14%)
WACC = 4.21875 + 3.50%
WACC = 7.72%

Therefore, appropriate discount rate assuming average risk is 7.72%.